# Improving Predictions of Convective Storm Wind Gusts through Statistical Post-Processing of Neural Weather Models

### v1

@@ 1 Introduction
Figure 1: Pangu-Weather forecasts for a day with convective conditions are clipped to Switzerland and post-processed using one of our different models. The best results are obtained with the CNN, which takes entire maps of features as input. Five different generalised extreme-value distributions are created, one for each cluster. This example shows predictions for June 2nd, 2021 at 17:00 with a lead time of 27 h. Accuracy varies with the prediction (e.g., forecast for Cluster 2 is less accurate than for Cluster 4).
@@
Severe winds can significantly damage buildings [klawa2003model, pinto2007changing, donat2011high, prahl2012applying, prettenthaler2012risk] and pose risks to human lives [schoen_climatology_2011, black_nontornadic_2010, ashley_spatial_2007]. Responding to such extreme events requires providing public authorities with accurate and easily understandable forecasts in time to implement prevention measures [he2021wind]. Effective strategies rely on timely warnings, from long-term to short-term, and accurate medium-range forecasts (2–10 days) have proven particularly efficient when combined with impact forecasts [schroeter_forecasting_2021].
@@
Meanwhile, machine learning-based models — especially Neural Weather Models (NWMs) [rasp2024weatherbench] — are emerging as a viable complement to traditional numerical weather prediction (NWP) models [ebert2023outlook] for short to medium-range weather forecasts. For example, GraphCast [lam_graphcast_2023], Pangu-Weather [bi_accurate_2023], FourCastNet [pathak_fourcastnet_2022], and AIFS [lang2024aifs] can issue high-quality, deterministic and global forecasts with a 0.25 spatial resolution for lead times up to seven days (ten for GraphCast). Once trained, these models can be run locally on a personal computer on a GPU in seconds. In contrast, physics-based NWP models solve non-linear partial differential equations using dedicated high-performance computing infrastructure. Although their outputs are generally made publicly available by national weather services and thus widely utilised, this complexity and cost prevent the general public from running them, hindering smaller weather prediction centres and hazard-issuing organisations.
@@
The use of NWMs for severe weather forecasting is rapidly evolving, from tropical cyclone tracking [bi_accurate_2023] to predicting local surface extremes [price2025probabilistic]. However, forecasting thunderstorms — highly localised, short-lived, and inherently uncertain events — presents unique challenges. Although NWMs have typical spatial and temporal resolutions that are reasonable to forecast mesoscale environments conducive to severe weather [feldmann2024convective], they currently lack the resolution and variables (e.g., precipitation) needed to directly detect individual convective storms, which often last only a few hours and cover areas of just tens of square kilometres. To address this, we propose post-processing the atmospheric environments predicted by NWMs to forecast regional distributions of wind gusts (they can be seen as the distributions of possible wind gusts for a “representative” site within the region) rather than deriving the tracks of individual storms from forecasted fields. Although hail often remains the primary source of thunderstorm damage [feldmann_hailstorms_2023], we here focus on wind gusts as they depend on more predictable factors such as mean wind, shear, and topography [sheridan_current_2018], while still posing significant risks to infrastructure and public safety.
@@
There are various methods to post-process NWP models to forecast station-wise wind gusts. For example, Friederichs et al. [friederichs_forecast_2012] used Vector Generalised Additive Models (VGAMs) to predict station-wise Generalised Extreme-value (GEV) distributions of wind gusts and Generalised Linear Models (VGLMs), Schulz et al. [schulz_machine_2022] reviewed different statistical post-processing techniques (multi-linear regression, machine learning, and neural networks) to post-process wind gust distributions, and Fontana et al. [fontana_convolutional_nodate] post-processed deterministic wind gusts forecasts via a Convolutional Neural Network (CNN) to predict distributions. As the considered wind gusts are maxima over 1 hour of the 1-s sustained wind speed (i.e., maxima over $3600$ observations), the GEV distribution, which is the asymptotic distribution of block maxima under appropriate conditions [[, e.g.,]]coles_introduction_2001, appears as a natural model for them [[, e.g., ]]friederichs_forecast_2012,schulz_machine_2022; their suitability in our context is expounded through diagnostic plots and Kolmogorov–Smirnov tests in the methods section. However, to our knowledge, these methods have not yet been applied to NWMs to predict regional distributions of meteorological variables.
@@
Here, we post-process a NWM to issue probabilistic forecasts of regional hourly wind gusts using a hierarchy of empirical models (statistical and machine learning-based). We selected the Pangu-Weather NWM for this study because it is currently the only established NWM that generates medium-range forecasts with an hourly temporal resolution. Our study focuses on Switzerland, allowing us to leverage high-quality wind gust observations over complex topography to assess thunderstorm-induced wind gust hazards. However, our framework can easily be adapted to other regions with station-based or gridded wind gust observations, enabling wind gust forecasts in areas where NWP models are not routinely run at the fine resolution needed to resolve extremes.
@@ 2 Results
Four post-processing models were trained on five years of Pangu-Weather forecasts (2016–2020) following a k-fold cross-validation scheme with each fold corresponding to one year in the training set. Each model predicts the location, scale, and shape parameters of five wind gust GEV distributions for each time step and each lead time, corresponding respectively to five clusters of stations from the Swiss Meteorological Network (SwissMetNet, [noauthor_automatic_2018]). The clusters are pre-computed based on measurement correlation between stations (see Section 4 for more details). Performing a forward sequential feature selection with a VGLM led us to post-process only four fields forecasted by the NWM, namely 10 m-wind ($u_{10m}$), 2 m-temperature ($T_{2m}$), geopotential at 500 hPa ($\phi_{500hPa}$), and wind at 850 hPa ($u_{850hPa}$) (see Section  4). The models considered were then evaluated on the year 2021, filtered to keep only dates when Switzerland experienced a thunderstorm to account for convective wind storm presence. The models include a VGAM based on the evgam [youngman_evgam_2020] R package, an Artificial Neural Network (ANN) trained on the bilinear interpolation of Pangu-Weather forecasts on SwissMetNet stations, and a CNN and a Vision Transformer (ViT) trained on Pangu-Weather outputs clipped to Switzerland (Figure 1). All Neural Networks (NNs) were implemented in Python with the jax library [bradbury_jax_2018].
Neural networks were trained using the Continuous Ranked Probability Score (CRPS) [friederichs_forecast_2012] as a loss function, and the VGAM was calibrated by maximum likelihood estimation. To compare the different models across datasets, another metric was introduced: the Continuous Ranked Probability Skill Score (CRPSS), defined as
@@
where “climatology” corresponds to the empirical distribution of wind gusts across stations and times for each cluster, obtained with all wind gust measurements from April to October for years 2016 to 2020, and “best possible” refers to the empirical distribution of wind gusts on each cluster at a specific time. As the models aim at fitting a distribution in a cluster of stations, the “best possible” CRPS cannot be outperformed. Therefore, CRPSS varies from negative values to 1 (best possible model), with 0 corresponding to climatology (no skill). To benchmark post-processing models against a state-of-the-art NWP system, we also evaluated the direct output of the ICON-CH2-EPS ensemble (referred to as ICON). For this, MeteoSwiss provided operational 10-m wind gust forecasts for the test year 2021.
Encouragingly, Fig. 2a shows that not only all post-processing models are skilful, but also outperform the ICON model, particularly at longer lead times. The gap between our best post-processing model (the CNN) and ICON is 0.2 for lead times under 32 hours but exceeds 0.3 for longer lead times. Pangu-Weather was initialized with ERA5 reanalysis, but previous research has shown that changing from reanalysis to operational analysis has little impact on the outcome of Pangu-Weather forecasts [feldmann2024convective], in line with NWMs’ ensembles being underdispersive compared to traditional ensemble methods [bi_accurate_2023]. We thus expect our results to be robust with regard to the choice of initial conditions. The similar sawtooth pattern for all models results from the use of different test sets at each lead time to accommodate the requirement of having a thunderstorm on the forecast date and the constraint of ICON being launched every 18 hours (see Section 4).
@@
To assess whether post-processing Pangu-Weather provides information for wind gust prediction, other models were trained as direct forecasts from the prediction dates, using the ERA5 dataset as initial conditions: regional distributions for time $t+\Delta t$ were forecasted from ERA5 data at time $t$. Panel (c) depicts the variation of CRPSS across lead times for all four ERA5-based models against the Pangu-based ones — for which the forecasts for time $t+\Delta t$ were obtained by post-processing a Pangu-Weather forecast for lead time $\Delta t$ computed from ERA5 data at time $t$.
@@
Figure 3 shows that, except with light wind gusts, the CNN is the best model. The VGAM experiences a sharper decrease in its performance for higher wind gusts than other models (69% decrease on average for the VGAM vs. 29% for the CNN, 42% for the ANN, and 42% for the ViT on moderate, strong, and extreme winds against light wind). Interestingly, all models perform better for extreme wind gusts than for moderate and strong ones. Extreme gusts are driven by specific conditions with high instability, and strong mid-level winds leading to fast-moving storms. These may be clearer to identify than intermediate cases producing less intense gusts (see Fig. 4c).
@@
Figure 4: Near-surface wind speed, near-surface temperature, 850 hPa wind and 500 hPa geopotential are used across models to predict wind gust distributions over each cluster — a task for which the CNN primarily focuses on local information. Panel (a) shows the mean absolute SHAP value and its min/max for each feature of the CNN over the five folds, based on a random sample of 50 dates for cluster 1 and for the prediction of the location parameter $\mu$ of wind gusts. Panel (b) shows the mean absolute SHAP value for $u_{10m}$, based on the same samples (black dots represent the stations of cluster 1). Panel (c) represents the VGAM partial dependence plot for the two main predictors for $\mu$ and $\sigma$, along with the univariate distributions of said predictors.
To gain a deeper understanding of our best post-processing models, we compute SHapley Additive exPlanations (SHAP, [lundberg2017unified]) using the game theory-based shap [noauthor_welcome_nodate] interpretability library in Python. Shapley values quantify how each input feature shifts a specific model prediction away from its expected (mean) value. By comparing the mean absolute SHAP value associated with each feature, we can assess their relative importance, including that of meteorological variables and the spatial importance of pixels within their fields.
@@
Figure 4c provides insight into the contribution of $u_{10m}$ and $u_{850hPa}$ to wind gust forecasts. The wind gusts’ location parameter increases with $u_{10m}$, with a coefficient of 1.4. This is significantly smaller than the theoretical wind gust factor for gusts measured over one second and mean wind measured over one hour, which has a theoretical value of 3.4 [noauthor_gust_nodate]. The parameter $\sigma$ also increases with $u_{850hPa}$, with a coefficient of 0.5.
@@ 4 Methods
We chose to post-process Pangu-Weather to take full advantage of the hourly time resolution of its outputs. Pangu-Weather is an NWM consisting of a 3D Earth-specific transformer trained on the ERA5 dataset, a 43-year meteorological reanalysis provided by the European European Centre for Medium-Range Weather Forecasts [bi_accurate_2023]. It provides forecasts of 4 surface variables and 5 upper air variables for 13 pressure levels (Figure 2) at a spatial resolution of 0.25° $\times$ 0.25°.
@@
We used Pangu-Weather to produce a training set of 5 years (2016 to 2020) and a test set of one year (2021, on which Pangu-Weather was not trained) clipped to Switzerland, each year containing the months April to October as they correspond to the convective season described in [feldmann_hailstorms_2023]. To balance resolution and cost, for each date, we produced forecasts with hourly lead times up to 1 day, 3-hourly lead times up to 2 days, and 12-hourly lead times up to 3 days.
In addition to Pangu-Weather’s standard variables (Figure 2), two storm-related variables were calculated with the Python library wrf-python [technologies_geoscience_2024]: the maximum CAPE in the lowest layers to assess the environmental convective instability following the method from [allen_severe_2011], and the 0–3 km storm-relative helicity (SRH) to assess tornadic instability and supercell presence [kerr_storm-relative_1996].
@@
Wind gusts were obtained through the SwissMetNet dataset [noauthor_automatic_2018], a network of 157 automatic stations disseminated over Switzerland providing measurements of different meteorological variables such as precipitation, temperature, air humidity, and wind every 10 minutes. Hourly measurements of wind gusts (in m s-1) — defined as the maximum wind speed over the preceding hour sustained during 1 second — were used as the target variable. The number of working SwissMetNet stations varies from one month to the other, as some stations were added or repaired between 2016 and 2021. We alleviated this issue by filtering the stations so that only those working through at least 70% of the dates in the dataset were kept. Finally, only the dates with complete station wind gust measurements data were kept.
@@
The storm tracks provided in the dataset used by [feldmann_hailstorms_2023] were used to evaluate the different models on severe events and to filter storm events in the training set. A supplementary filter on the storms was then added to keep only the ones detected by a station, as the dataset also contained storms detected by Swiss radar measurements in neighbouring countries.
@@
The ICON-CH2-EPS data used in this study are part of a set of reforecasts covering April to October 2021 with $18$ h update frequency and lead times ranging from 0 to 72 h. From the 21-member ensemble forecast, station-wise 10 m wind gust values were obtained by finding the nearest pixel to the SwissMetNet stations on the 2.2 km grid.
@@
The Continuous Ranked Probability Score (CRPS) was chosen as a metric to train and evaluate the different models. As a proper scoring rule validating all three of Murphy’s forecast goodness types [murphy_what_1993], it is widely used to assert weather model performance ([bi_accurate_2023], [schulz_machine_2022], [bremnes_evaluation_2024]). Moreover, CRPS-based training has already provided the best forecasts for high-speed wind gusts in other studies [friederichs_forecast_2012].
@@
Models
@@
Interestingly, even though no direct spatial indicator was given as an input for the clustering (i.e., no indication of position), the obtained clusters are spatially coherent: they do not intersect, present climatic similarities, and approximately align with the regions used by MeteoSwiss to define the thresholds for extreme wind hazard warnings [MeteoSwiss2021]. We observe a clear separation between various homogeneous climate zones: the Alps (red for western Alps and black for eastern Alps), southern plains (orange), western plains and Jura (yellow), and northern plains (purple). The climatic stationarity observed in the obtained zones supports our assumption of constant values for $\mu$, $\sigma$, and $\xi$ across each region.
@@
for values of $x$ such that $01+\xi(x-\mu)/\sigma>$ 0. When $\xi=0$, the distribution corresponds to the limit of (3) as $\xi\to 0$.
As wind gusts are defined as the maximum wind speed over a certain time period, modelling them using GEV distributions is often suitable (e.g, [pinheiro_comparative_2016], [cheng_generalized_2002], [friederichs_probabilistic_2009]). Moreover, a notable advantage of GEV distributions in training machine learning models is the availability of a closed-form expression for the CRPS [friederichs_forecast_2012] that can be implemented as a loss function.
@@
Figure 9 shows the results of the feature selections: the first four variables are wind speed at 10 m ($u_{10m}$), temperature at 2 m ($T_{2m}$), geopotential at 500 hPa ($\phi_{500hPa}$), and wind at 850 hPa ($u_{850hPa}$). It is unsurprising: wind at 10 m gives a mean value that wind gusts exceed, temperature and geopotential give information about small-scale and large-scale pressure systems, and wind at 850 hPa can be interpreted as the influence of lower-atmospheric winds on surface wind gusts. More surprising is that none of the storm-related variables that were computed from Pangu-Weather outputs such as CAPE, the Level of Free Convection (LFC), or SRH, were selected in the first four variables. This could be linked to these variables being good predictors of a storm presence, but not necessarily of the wind gusts a storm induces, which are more related to Downward Convective Available Potential Energy (DCAPE), storm motion and entrainment of dry air [markowski2010hazards].
@@
Following [friederichs_forecast_2012], we first train Vector Generalised Linear and Additive Models (VGLMs and VGAMs) to predict the parameters of wind gust GEV distributions from Pangu-Weather output.
@@
Here, the logarithm ensures the scale parameter $\sigma$ remains positive. The shape parameter $\xi$ is kept constant—i.e., independent of the features $\mathbf{x}$—to prevent numerical instability [coles_introduction_2001]. To further ensure stability, $\xi$ is constrained to the interval $(-0.5,1)$ using the function $s$, defined as the inverse of the sigmoid function.
@@
Following [friederichs_forecast_2012], we first trained station-wise VGAMs and VGLMs. In total, 165 models were fitted: one for each of the 33 lead times and each of the 5 clusters. The models would thus predict different $\mu$ and $\sigma$ for each station but fixed $f_{.}^{(j)}$ for each cluster and lead time. The features were in this case the interpolation of Pangu-Weather outputs on each station.
Aiming to fit one distribution per cluster, the input covariates were replaced by the mean value of each feature on the cluster (average of the interpolated values at the stations). Interestingly, the cluster-wise VGAM eventually outperformed the station-wise VGAM (CRPSS of 0.41 against 0.26). It could be explained by the lack of small-scale spatial accuracy in Pangu-Weather outputs that a spatial aggregation eludes. In the results section, only the cluster-wise VGAM is presented as it gave the best results.

### v3

@@ 1 Introduction
Figure 1: Pangu-Weather forecasts for a day with convective conditions are clipped to Switzerland and post-processed using one of our different models. The best results are obtained with the CNN, which takes entire maps of features as input. Five different generalised extreme-value distributions are created, one for each cluster. This example shows predictions for June 2nd, 2021 at 17:$0$ with a lead time of 27 h. Accuracy varies with the prediction (e.g., forecast for Cluster 2 is less accurate than for Cluster 4).
@@
Severe winds can significantly damage buildings [1, 2, 3, 4, 5] and pose risks to human lives [6, 7, 8]. Responding to such extreme events requires providing public authorities with accurate and easily understandable forecasts in time to implement prevention measures [9]. Effective strategies rely on timely warnings, from long-term to short-term, and accurate medium-range forecasts (2–10 days) have proven particularly efficient when combined with impact forecasts [10].
@@
Meanwhile, machine learning-based models — especially Neural Weather Models (NWMs) [11] — are emerging as a viable complement to traditional numerical weather prediction (NWP) models [12] for short to medium-range weather forecasts. For example, GraphCast [13], Pangu-Weather [14], FourCastNet [15], and AIFS [16] can issue high-quality, deterministic and global forecasts with a 0.25 spatial resolution for medium range lead times (up to seven days). Once trained, these models can be run locally on a personal computer on a GPU in seconds. In contrast, physics-based NWP models solve non-linear partial differential equations using dedicated high-performance computing infrastructure. Although their outputs are generally made publicly available by national weather services and thus widely utilised, this complexity and cost prevent the general public from running them, hindering smaller weather prediction centres and hazard-issuing organisations.
@@
The use of NWMs for severe weather forecasting is rapidly evolving, from tropical cyclone tracking [14] to predicting local surface extremes [17]. However, forecasting thunderstorms — highly localised, short-lived, and inherently uncertain events — presents unique challenges. Although NWMs have typical spatial and temporal resolutions that are reasonable to forecast mesoscale environments conducive to severe weather [18], they currently lack the resolution and variables (e.g., precipitation) needed to directly detect individual convective storms, which often last only a few hours and cover areas of just tens of square kilometres. To address this, we propose post-processing the atmospheric environments predicted by NWMs to forecast regional distributions of wind gusts (they can be seen as the distributions of possible wind gusts for a “representative” site within the region) rather than deriving the tracks of individual storms from forecasted fields. Although hail often remains the primary source of thunderstorm damage [19], we here focus on wind gusts as they depend on more predictable factors such as mean wind, shear, and topography [20], while still posing significant risks to infrastructure and public safety.
@@
There are various methods to post-process NWP models to forecast station-wise wind gusts. For example, Friederichs et al. [21] used Vector Generalised Additive Models (VGAMs) to predict station-wise Generalised Extreme-value (GEV) distributions of wind gusts and Vector Generalised Linear Models (VGLMs), Schulz et al. [22] reviewed different statistical post-processing techniques (multi-linear regression, machine learning, and neural networks) to post-process wind gust distributions, and Fontana et al. [23] post-processed deterministic wind gusts forecasts via a Convolutional Neural Network (CNN) to predict distributions. As the considered wind gusts are maxima over 1 hour of the 1-s sustained wind speed (i.e., maxima over $3600$ observations), the GEV distribution, which is the asymptotic distribution of block maxima under appropriate conditions [[, e.g.,]]coles_introduction_2001, appears as a natural model for them [[, e.g., ]]friederichs_forecast_2012,schulz_machine_2022; their suitability in our context is expounded through diagnostic plots and Kolmogorov–Smirnov tests in the methods section. However, to our knowledge, these methods have not yet been applied to NWMs to predict regional distributions of meteorological variables.
@@
Here, we post-process a NWM to issue probabilistic forecasts of regional hourly wind gusts using a hierarchy of empirical models (statistical and machine learning-based). We selected the Pangu-Weather NWM for this study because it is currently the only established NWM that generates medium-range forecasts with an hourly temporal resolution. Our study focuses on Switzerland, allowing us to leverage high-quality wind gust observations over complex topography to assess wind gust hazards under convectively active conditions. However, our framework can easily be adapted to other regions with station-based or gridded wind gust observations, enabling wind gust forecasts in areas where NWP models are not routinely run at the fine resolution needed to resolve extremes.
@@ 2 Results
Four post-processing models were trained on five years of Pangu-Weather forecasts (2016–2020) following a k-fold cross-validation scheme with each fold corresponding to one year in the training set. Each model predicts the location, scale, and shape parameters of five wind gust GEV distributions for each time step and each lead time, corresponding respectively to five clusters of stations from the Swiss Meteorological Network (SwissMetNet, [25]). The clusters are pre-computed based on measurement correlation between stations (see Section 4 for more details). Performing a forward sequential feature selection with a VGLM led us to post-process only four fields forecasted by the NWM, namely 10 m-wind ($u_{10m}$), 2 m-temperature ($T_{2m}$), geopotential at 500 hPa ($\phi_{500hPa}$), and wind at 850 hPa ($u_{850hPa}$) (see Section  4). The models considered were then evaluated on the year 2021, filtered to keep only dates when Switzerland experienced a thunderstorm to account for convective activity. The models include a VGAM based on the evgam [26] R package, an Artificial Neural Network (ANN) trained on the bilinear interpolation of Pangu-Weather forecasts on SwissMetNet stations, and a CNN and a Vision Transformer (ViT) trained on Pangu-Weather outputs clipped to Switzerland (Figure 1). All Neural Networks (NNs) were implemented in Python with the jax library [27].
Neural networks were trained using the Continuous Ranked Probability Score (CRPS) [21] as a loss function, and the VGAM was calibrated by maximum likelihood estimation. To compare the different models across datasets, another metric was introduced: the Continuous Ranked Probability Skill Score (CRPSS), defined as
@@
where “climatology” corresponds to the empirical distribution of wind gusts across stations and times for each cluster, obtained with all wind gust measurements from April to October for years 2016 to 2020, and “best possible” refers to the empirical distribution of wind gusts on each cluster at a specific time. As the models aim at fitting a distribution in a cluster of stations, the “best possible” CRPS cannot be outperformed. Therefore, CRPSS varies from negative values to 1 (best possible model), with 0 corresponding to climatology (no skill). To benchmark post-processing models against a state-of-the-art NWP system, we also evaluated the direct output of the ICON-CH2-EPS ensemble (referred to as ICON). For this, MeteoSwiss provided operational 10-m wind gust forecasts for the test year 2021 (see Supplementary Figure 8, Supplementary Information section 5 for more details about how the test year was chosen).
Encouragingly, Figure 2a shows that not only all post-processing models are skilful, but also outperform the ICON model, particularly at longer lead times. The gap between our best post-processing model (the CNN) and ICON is 0.2 for lead times under 32 hours but exceeds 0.3 for longer lead times. Pangu-Weather was initialized with ERA5 reanalysis, but previous research has shown that changing from ERA5 reanalysis to IFS operational analysis has little impact on the outcome of Pangu-Weather forecasts [18], in line with NWMs’ ensembles being underdispersive compared to traditional ensemble methods [14]. We thus expect our results to be robust with regard to the choice of initial conditions. The similar sawtooth pattern for all models results from the use of different test sets at each lead time to accommodate the requirement of having a thunderstorm on the forecast date and the constraint of ICON being launched every 18 hours (see Section 4).
@@
To assess whether post-processing Pangu-Weather provides information for wind gust prediction or not, as NNs could be powerful enough to issue gust forecast from raw observations, other models with the same architecture were trained as simple forecasting baselines from the prediction dates with ERA5 dataset as initial conditions. Regional distributions for time $t+\Delta t$ were forecasted from ERA5 data at time $t$. Panel (c) depicts the variation of CRPSS across lead times for all four ERA5-based models against the Pangu-based ones — for which the forecasts for time $t+\Delta t$ were obtained by post-processing a Pangu-Weather forecast for lead time $\Delta t$ computed from ERA5 data at time $t$.
@@
Figure 3 shows that, except with light wind gusts, the CNN is the best model. The VGAM experiences a sharper decrease in its performance for higher wind gusts than other models (69% decrease on average for the VGAM vs. 29% for the CNN, 42% for the ANN, and 42% for the ViT on moderate, strong, and extreme winds against light wind). Interestingly, all models perform better for extreme wind gusts than for moderate and strong ones (see Supplementary Figures 3 to 5, SI section S3 for investigation on sensitivity to filtering). Extreme gusts are driven by specific conditions with high instability, and strong mid-level winds leading to fast-moving storms. These may be clearer to identify than intermediate cases producing less intense gusts (see Figure 4c).
@@
Figure 4: Near-surface wind speed, near-surface temperature, 850 hPa wind and 500 hPa geopotential are used across models to predict wind gust distributions over each cluster — a task for which the CNN primarily focuses on local information. Panel (a) shows the mean absolute SHAP value and its min/max for each feature of the CNN over the five folds, based on a random sample of 50 dates for cluster 1 and for the prediction of the location parameter $\mu$ of wind gusts. Panel (b) shows the mean absolute SHAP value for $u_{10m}$, based on the same samples (black dots represent the stations of cluster 1). Panel (c) represents the VGAM partial dependence plot for the two main predictors for $\mu$ and $\sigma$, along with the univariate distributions of said predictors. It shows how an increase in 10m wind implies more important wind gust, and how an increase in 850hPa winds implies more dispersion of wind gust distribution.
To gain a deeper understanding of our best post-processing models, we compute SHapley Additive exPlanations (SHAP, [28]) using the game theory-based shap [29] interpretability library in Python. Shapley values quantify how each input feature shifts a specific model prediction away from its expected (mean) value. By comparing the mean absolute SHAP value associated with each feature, we can assess their relative importance, including that of meteorological variables and the spatial importance of pixels within their fields.
@@
Figure 4c provides insight into the contribution of $u_{10m}$ and $u_{850hPa}$ to wind gust forecasts. The wind gusts’ location parameter increases with $u_{10m}$, with a coefficient of 1.4. This is significantly smaller than the theoretical wind gust factor for gusts measured over one second and mean wind measured over one hour, which has a theoretical value of 3.4 [30]. The parameter $\sigma$ also increases with $u_{850hPa}$, with a coefficient of 0.5.
@@ 4 Methods
We chose to post-process Pangu-Weather to take full advantage of the hourly time resolution of its outputs. Pangu-Weather is an NWM consisting of a 3D Earth-specific transformer trained on the ERA5 dataset, a 43-year meteorological reanalysis provided by the European European Centre for Medium-Range Weather Forecasts [14]. It provides forecasts of 4 surface variables and 5 upper air variables for 13 pressure levels (Table 2) at a spatial resolution of 0.25° $\times$ 0.25°. Pangu is a succession of four models which were each trained for a specific lead time (1h, 3h 6h, and 24h). To produce a forecast, a minimal number of Pangu submodels is used, from the longest to the shortest lead time.
@@
We used Pangu-Weather to produce a training set of 5 years (2016 to 2020) and a test set of one year (2021, on which Pangu-Weather was not trained) clipped to Switzerland, each year containing the months April to October as they correspond to the convective season described in [19]. To balance resolution and cost, for each date, we produced forecasts with hourly lead times up to 1 day, 3-hourly lead times up to 2 days, and 12-hourly lead times up to 3 days.
In addition to Pangu-Weather’s standard variables (Table 2), two storm-related variables were calculated with the Python library wrf-python [31]: the maximum CAPE in the lowest layers to assess the environmental convective instability following the method from [32], and the 0–3 km above ground storm-relative helicity (SRH) to assess tornadic instability and supercell presence [33].
@@
Wind gusts were obtained through the SwissMetNet dataset [25], a network of 157 automatic stations disseminated over Switzerland providing measurements of different meteorological variables such as precipitation, temperature, air humidity, and wind every 10 minutes. Hourly measurements of wind gusts (in m s-1) — defined as the maximum wind speed over the preceding hour sustained during 1 second — were used as the target variable. The number of working SwissMetNet stations varies from one month to the other, as some stations were added or repaired between 2016 and 2021. We alleviated this issue by filtering the stations so that only those working through at least 70% of the dates in the dataset were kept. Finally, only the dates with complete station wind gust measurements data were kept.
@@
The storm tracks provided in the dataset used by [19] were used to evaluate the different models on severe events and to filter storm events in the training set. A supplementary filter on the storms was then added to keep only the ones detected by a station, as the dataset also contained storms detected by Swiss radar measurements in neighbouring countries.
@@
As ICON-CH2-EPS did not run yet in 2021, the ICON data used in this study are part of a set of reforecasts covering April to October 2021 with a reduced update frequency (every $18$ h) and lead times ranging from 0 to 72 h. From the 21-member ensemble forecast, station-wise 10 m wind gust values - computed as hourly maximum of convective and turbulent components of wind gust - were obtained by finding the nearest pixel to the SwissMetNet stations on the 2.2 km grid.
@@
The Continuous Ranked Probability Score (CRPS) was chosen as a metric to train and evaluate the different models. As a proper scoring rule validating all three of Murphy’s forecast goodness types [35], it is widely used to assert weather model performance ([14], [22], [36]). Moreover, CRPS-based training has already provided the best forecasts for high-speed wind gusts in other studies [21].
@@
SHAP values
For each model $m$, the mean absolute SHAP value of a variable $u$ is defined by :
$\begin{split}|\overline{\text{SHAP}(u_{m})}|\ =\\ \sum_{(x,y,\Delta t,T)\in S}&|\text{SHAP}\big{(}\ u_{m}(x,y,\Delta t,t)\ \big{)}|\end{split}$ (3)
where $u$ is the predicted variable, $x$ is the longitude, $y$ is the latitude, $\Delta t$ is the lead time, $T$ is the targeted prediction date, and $S$ is a sample of prediction date with all lead times, all longitude, all latitude.
Figure 4 (a) shows the min-max of $|\overline{\text{SHAP}(u_{m})}|$ for $m\in[1,2,3,4,5]$ being the five CNN parametrisation from the k-fold training.
For Figure 4 (b), at each point, the SHAP value is given by :
$\begin{split}|\overline{\text{SHAP}\big{(}u(x,y\big{)}}|\ =\\ \sum_{(\Delta t,T,m)\in R}&|\text{SHAP}\big{(}\ u_{m}(\Delta t,t)\ \big{)}|\end{split}$ (4)
where $u$ is the predicted variable, $x$ is the longitude, $y$ is the latitude, $\Delta t$ is the lead time, $T$ is the targeted prediction date, and $R$ is a sample of prediction date with all lead times, all models.
@@
Interestingly, even though no direct spatial indicator was given as an input for the clustering (i.e., no indication of position), the obtained clusters are spatially coherent: they do not intersect, present climatic similarities, and approximately align with the regions used by MeteoSwiss to define the thresholds for extreme wind hazard warnings [37]. Figure 7 reveals a clear separation between various homogeneous climate zones: the Alps (red for western Alps and black for eastern Alps), southern plains (orange), western plains and Jura (yellow), and northern plains (purple). The climatic stationarity observed in the obtained zones supports our assumption of constant values for $\mu$, $\sigma$, and $\xi$ across each region.
@@
for values of $x$ such that $1+\xi(x-\mu)/\sigma>0$. When $\xi=0$, the distribution corresponds to the limit of (5) as $\xi\to 0$.
As wind gusts are defined as the maximum wind speed over a certain time period, modelling them using GEV distributions is often suitable (e.g, [38], [39], [40]). Moreover, a notable advantage of GEV distributions in training machine learning models is the availability of a closed-form expression for the CRPS [21] that can be implemented directly as a loss function:
For $\xi\neq 0$:
$\begin{split}C&RPS(F_{(\mu,\sigma,\xi)},y)=(\mu-y-\frac{\sigma}{\xi})(1-2F_{\mu,\sigma,\xi})\\ &-\frac{\sigma}{\xi}(2^{\xi}\Gamma(1-\xi)-2\Gamma_{l}(1-\xi,-\log F_{\xi,\sigma,\mu}(y)))\end{split}$ (6)
For $\xi=0$:
$\begin{split}CRPS(F_{(\mu,\sigma,0)},y)&=\mu-y+\sigma(\gamma-\log 2)\\ &-2\sigma Ei(\log F_{\mu,\sigma,0}(y))\end{split}$ (7)
@@
Figure 9 shows the results of the feature selections: the first four variables are wind speed at 10 m ($u_{10m}$), temperature at 2 m ($T_{2m}$), geopotential at 500 hPa ($\phi_{500hPa}$), and wind at 850 hPa ($u_{850hPa}$). It is unsurprising: wind at 10 m gives a mean value that wind gusts exceed, temperature and geopotential give information about small-scale and large-scale pressure systems, and wind at 850 hPa can be interpreted as the influence of lower-atmospheric winds on surface wind gusts. More surprising is that none of the storm-related variables that were computed from Pangu-Weather outputs such as CAPE, the Level of Free Convection (LFC), or SRH, were selected in the first four variables. This could be linked to these variables being good predictors of a storm presence, but not necessarily of the wind gusts a storm induces, which are more related to Downward Convective Available Potential Energy (DCAPE), storm motion and entrainment of dry air [41].
@@
Following [21], we first train Vector Generalised Linear and Additive Models (VGLMs and VGAMs) to predict the parameters of wind gust GEV distributions from Pangu-Weather output.
@@
Here, the logarithm ensures the scale parameter $\sigma$ remains positive. The shape parameter $\xi$ is kept constant—i.e., independent of the features $\mathbf{x}$—to prevent numerical instability [24]. To further ensure stability, $\xi$ is constrained to the interval (0.5,1)(-0.5,1) using the function $s$, defined as the inverse of the sigmoid function.
@@
Following [21], we first trained station-wise VGAMs and VGLMs. In total, 165 models were fitted: one for each of the 33 lead times and each of the 5 clusters. The models would thus predict different $\mu$ and $\sigma$ for each station but fixed $f_{.}^{(j)}$ for each cluster and lead time. The features were in this case the interpolation of Pangu-Weather outputs on each station.
Aiming to fit one distribution per cluster, the input covariates were replaced by the mean value of each feature on the cluster (average of the interpolated values at the stations). Interestingly, the cluster-wise VGAM outperformed the station-wise VGAM (CRPSS of 0.41 against 0.26). It could be explained by the limited spatial resolution of Pangu-Weather, whose global-scale design lacks the small-scale spatial accuracy needed to benefit fully from station-specific modeling, and the unpredictability of individual stations, making individual forecasts less precise - even when forecast is a distribution. Additionally, using cluster-averaged features allowed us to avoid the complexity and computational cost of fitting 5 VGAM with approximately 120 covariates (4 variables across about 30 stations for each cluster). In the results section, only the cluster-wise VGAM is presented as it gave the best results. The same reasons ultimately lead us to prefer cluster-wise fits rather than station-based ones for the NNs to tackle Pangu inherent coarse resolution (being a global model), individual station unpredictability, and computational training and running cost.